# Further author information: (Send correspondence to Yuankai Huo)

### v1

@@ Abstract
Accurate lung cancer risk prediction remains challenging due to substantial variability across patient populations and clinical settings—no single model performs best for all cohorts. To address this, we propose personalized lung cancer risk prediction agent that dynamically selects the most appropriate model for each patient by combining cohort-specific knowledge with modern retrieval and reasoning techniques. Given a patient’s CT scan and structured metadata—including demographic, clinical, and nodule-level features—the agent first performs cohort retrieval using FAISS-based similarity search across nine diverse real-world cohorts to identify the most relevant patient population from a multi-institutional database. Second, a Large Language Model (LLM) is prompted with the retrieved cohort and its associated performance metrics to recommend the optimal prediction algorithm from a pool of eight representative models, including classical linear risk models (e.g., Mayo, Brock), temporally-aware models (e.g., TDVIT, DLSTM), and multi-modal computer vision based approaches (e.g., Liao, Sybil, DLS, DLI). This two-stage agent pipeline—retrieval via FAISS and reasoning via LLM—enables dynamic, cohort-aware risk prediction personalized to each patient’s profile. Building on this architecture, the agent supports flexible and cohort-driven model selection across diverse clinical populations, offering a practical path toward individualized risk assessment in real-world lung cancer screening.
@@
To address this gap, we propose Cohort-Aware Agents, a retrieval-augmented framework for individualized lung cancer risk prediction. Rather than applying a fixed model across all patients, our approach automatically selects the most appropriate prediction algorithm by combining cohort retrieval with LLM-based reasoning. The agent follows a two-stage architecture inspired by clinical decision-making. In the first stage, it retrieves the most relevant reference cohort by comparing a patient’s CT scan and structured clinical metadata—including demographic [41], clinical, and nodule-level features—with a large, multi-institutional database using FAISS-based [42] similarity search. This allows the agent to match each patient to a representative clinical subpopulation, capturing subtle patterns in data distribution. In the second stage, a LLM is prompted with the retrieved cohort’s characteristics and associated model performance summaries to identify the most suitable prediction algorithm. Rather than applying a fixed model, our agent dynamically selects from eight risk models, including classical statistical scores (e.g., Mayo [7], Brock [8]), temporally-aware models (e.g., TDVIT [9], DLSTM [10]), and computer vision based approaches (e.g., Liao [43], Sybil [44], DLS [12], DLI [11]). Figure 1 illustrates this overall framework by contrasting our cohort-aware agent with a naive RAG baseline. Our cohort-aware agent enables personalized, context-aware model selection that adapts to cohort-specific characteristics, providing a practical and generalizable solution for individualized lung cancer risk assessment. In summary, our agent makes the following contributions:
@@ 3.1 Cohort Overview
Our reference database comprises nine real-world lung cancer cohorts collected from diverse institutions and clinical contexts, each providing structured metadata, chest CT imaging, and diagnostic labels. The VLSP cohort comes from the Vanderbilt Lung Screening Program and represents a low-dose CT screening population. The LIVUSPN [47] cohort includes individuals enrolled in a longitudinal imaging and biomarker study, covering both high-risk and screening-detected cases. From the Multi-Center Lung (MCL) project, we incorporate four institution-specific cohorts: MCL-VUMC (Vanderbilt University Medical Center), MCL-UPMC (University of Pittsburgh Medical Center), MCL-UCD (University of California, Davis), and MCL-DECAMP [48] (from the DECAMP consortium), representing a mixture of screening, incidental, and diagnostic settings. The BRONCH cohort includes patients undergoing diagnostic bronchoscopy and is enriched for pathologically confirmed malignancies. We also include two subsets from the National Lung Screening Trial (NLST) [49]: NLST-test [50], a held-out split of the original trial, and NLST-test-nodule, a curated subset containing patients with confirmed pulmonary nodules.
All cohorts provide structured metadata suitable for similarity-based cohort retrieval, including features such as age, gender, BMI, smoking status, and scan date. In addition, each cohort includes nodule-level imaging features derived from chest CTs, along with binary cancer outcome labels confirmed through pathology or long-term follow-up. This consistency across cohorts enables robust construction of patient representations and supports cohort-aware retrieval and model selection within our framework.
@@ 3.2 Model Pool
Our agent selects from a pool of eight lung cancer risk prediction models, representing a diverse set of modeling paradigms and clinical assumptions. These models are grouped into three broad categories:
Linear risk models. The Mayo [7] and Brock [8, 7] models are classical statistical scores based on logistic regression, widely used in clinical practice. They estimate malignancy risk using variables such as nodule size, patient age, smoking status, and spiculation. These models offer interpretability and strong baseline performance but are limited by their handcrafted feature [51, 52, 53] sets and lack of imaging representation.
Temporally-aware models. TDVIT [9] and DLSTM [10] incorporate temporal information from sequential scans to model nodule evolution over time. TDVIT uses a transformer architecture with scan-position embeddings, while DLSTM leverages recurrent structures to track longitudinal changes. These models are applicable only to cohorts with multiple timepoints available [54].
Computer vision based models. Liao [43], Sybil [44], DeepLungScreening (DLS) [12], and DeepLungIPN (DLI) [11] utilize CT imaging, often in combination with structured metadata, to learn high-dimensional predictive features. Liao and Sybil focus on nodule- and whole-lung-based representations, respectively. DLS is optimized for screening populations, while DLI is designed for incidental and diagnostic settings and requires additional metadata fields [55]. Because of their differing input requirements, not all models are applied to all cohorts.
Model selection is performed adaptively by the agent based on availability of required inputs and historical performance on retrieved cohorts.
@@ 4.1 Cohort-Aware Retrieval Performance
| bronch | mcl_vumc | mcl_upmc | mcl_decamp | mcl_ucd | vlsp | livuspn | nlst_test_nodule | nlst_test | correct / total (acc) | |

### v2

@@ Abstract
Accurate lung cancer risk prediction remains challenging due to substantial variability across patient populations and clinical settings—no single model performs best for all cohorts. To address this, we propose personalized lung cancer risk prediction agent that dynamically selects the most appropriate model for each patient by combining cohort-specific knowledge with modern retrieval and reasoning techniques. Given a patient’s CT scan and structured metadata—including demographic, clinical, and nodule-level features—the agent first performs cohort retrieval using FAISS-based similarity search across nine diverse real-world cohorts to identify the most relevant patient population from a multi-institutional database. Second, a Large Language Model (LLM) is prompted with the retrieved cohort and its associated performance metrics to recommend the optimal prediction algorithm from a pool of eight representative models, including classical linear risk models (e.g., Mayo, Brock), temporally-aware models (e.g., TD-VIT, DLSTM), and multi-modal computer vision based approaches (e.g., Liao, Sybil, DLS, DLI). This two-stage agent pipeline—retrieval via FAISS and reasoning via LLM—enables dynamic, cohort-aware risk prediction personalized to each patient’s profile. Building on this architecture, the agent supports flexible and cohort-driven model selection across diverse clinical populations, offering a practical path toward individualized risk assessment in real-world lung cancer screening.
@@
To address this gap, we propose Cohort-Aware Agents, a retrieval-augmented framework for individualized lung cancer risk prediction. Rather than applying a fixed model across all patients, our approach automatically selects the most appropriate prediction algorithm by combining cohort retrieval with LLM-based reasoning. The agent follows a two-stage architecture inspired by clinical decision-making. In the first stage, it retrieves the most relevant reference cohort by comparing a patient’s CT scan and structured clinical metadata—including demographic [41], clinical, and nodule-level features—with a large, multi-institutional database using FAISS-based [42] similarity search. This allows the agent to match each patient to a representative clinical subpopulation, capturing subtle patterns in data distribution. In the second stage, a LLM is prompted with the retrieved cohort’s characteristics and associated model performance summaries to identify the most suitable prediction algorithm. Rather than applying a fixed model, our agent dynamically selects from eight risk models, including classical statistical scores (e.g., Mayo [7], Brock [8]), temporally-aware models (e.g., TD-VIT [9], DLSTM [10]), and computer vision based approaches (e.g., Liao [43], Sybil [44], DLS [12], DLI [11]). Figure 1 illustrates this overall framework by contrasting our cohort-aware agent with a naive RAG baseline. Our cohort-aware agent enables personalized, context-aware model selection that adapts to cohort-specific characteristics, providing a practical and generalizable solution for individualized lung cancer risk assessment. In summary, our agent makes the following contributions:
@@ 3.1 Cohort Overview
Our reference database comprises nine real-world lung cancer cohorts collected across institutions and clinical contexts, each with structured metadata, chest CT imaging, and diagnostic labels. The VLSP cohort comes from the Vanderbilt Lung Screening Program and represents a low-dose CT screening population. The LI-VUMC [47] cohort is a longitudinal imaging and biomarker cohort at VUMC, covering both high-risk and screening-detected cases. From the Multi-Center Lung (MCL) project, we incorporate four institution-specific cohorts: MCL-VUMC (Vanderbilt University Medical Center), MCL-UPMC (University of Pittsburgh Medical Center), MCL-UCD (University of Colorado Denver), and MCL-DECAMP [48] (DECAMP consortium), representing prospectively enrolled incidental nodule populations across sites. The BRONCH cohort includes patients undergoing diagnostic bronchoscopy with pathology-adjudicated outcomes. We also include two subsets from the National Lung Screening Trial (NLST) [49]: NLST-test [50], a held-out split aligned with the original trial’s test set, and NLST-test-nodule, a subset of NLST-test with at least one nodule meeting the NLST positive definition.
All cohorts provide structured metadata suitable for similarity-based cohort retrieval (e.g., age, gender, BMI, smoking status, scan date). When nodule-level descriptors are unavailable, we rely on image-level features and cohort-level metadata; outcome labels follow each dataset’s ground-truth protocol (pathology, clinical coding, or longitudinal follow-up). This consistency enables robust patient representations and supports cohort-aware retrieval and model selection within our framework.
@@ 3.2 Model Pool
Our agent selects from a pool of eight lung cancer risk prediction models spanning diverse modeling paradigms and clinical assumptions. We group them into three categories:
Linear risk models. The Mayo [8] and Brock [7] models are logistic-regression scores widely used in practice, leveraging variables such as nodule size, age, smoking status, and spiculation. They offer interpretability and solid baselines but are constrained by hand-crafted feature sets [51, 52, 53] and the absence of learned imaging representations.
Temporally-aware models. TD-ViT [9] and DLSTM [10] incorporate longitudinal CT to capture nodule evolution across time. TD-ViT encodes continuous time/scan position into a transformer, while DLSTM uses recurrent gating to model temporal dynamics. These models are applicable only when multiple timepoints are available [54].
Computer vision-based models. Liao [43], Sybil [44], DeepLungScreening (DLS) [12], and DeepLungIPN (DLI) [11] learn high-dimensional predictors from CT, optionally fused with structured metadata. Sybil operates directly at the scan level without manual nodule annotations; DLS targets screening populations and integrates CT with clinical data elements; DLI targets indeterminate/incidental nodules with CT plus clinical variables [55]. Because input requirements differ, not all models are applied to all cohorts.
Model selection is performed adaptively by the agent, conditioned on available inputs and historical performance on retrieved cohorts.
@@ 4.1 Cohort-Aware Retrieval Performance
| BRONCH | MCL_VUMC | MCL_UPMC | MCL_DECAMP | MCL_UCD | VLSP | LI-VUMC | NLST_test_nodule | NLST_test | correct / total (acc) | |